Leary and O’Donnell, a local CPA firm, has been asked to bid on a contract to perform audits for three counties in its home state. Because existing staff are fully scheduled, if the firm is awarded the contract, it must hire one new staff member at a salary of $59,000 to handle the additional workload. The managing partner is convinced that obtaining the contract will lead to additional new clients from the respective counties. Expected new work (excluding the three counties) is 820 hours at an average billing rate of $90 per hour. Other information follows about the firm’s current annual revenues and costs:

Firm volume in hours (normal) 30,750
Fixed costs $ 575,000
Variable cost $ 35 /hr

Should the firm win the contract, the audits of the three counties will require 880 hours of expected work.

Required:

1. If the managing partner’s expectations are correct, what is the lowest bid the firm can submit and still expect to increase annual net income? What would be the hourly billing rate for the county audit jobs just to break even on all the new business? (Round "Average billing rate" answer to 2 decimal places.)

2. If the contract is obtained at a price of $44,700, what is the minimum number of hours of new business in addition to the county work that must be obtained for the firm to break even on total new business? What is the margin of safety (MOS) regarding the county audit job proposal?
Answer #1

Answer 1.

Calculating lowest bid:

Additional cost = 59000

Variable cost = hours * rate = 880 * 35 = 30800

Total cost = 59000 + 30800 = 89800

Additional contribution = 820 * [90-35] = 45100

Lowest bid = total cost - additional contribution = 89800 - 45100 = 44700

Average billing rate = lowest bid / no of hours = 44700 / 880 = 50.80

Answer 2.

MOS = lowest bid = 44700

minimum no of hours as in question is 820 hours